# EVM-Fusion: An Explainable Vision Mamba Architecture with Neural Algorithmic Fusion

### v3

@@ Abstract
Medical image classification is critical for clinical decision-making, yet demands for accuracy, interpretability, and generalizability remain challenging. This paper introduces EVM-Fusion, an Explainable Vision Mamba architecture featuring a novel Neural Algorithmic Fusion (NAF) mechanism for multi-organ medical image classification. EVM-Fusion leverages a multipath design, where DenseNet and U-Net based pathways, enhanced by Vision Mamba (Vim) modules, operate in parallel with a traditional feature pathway. These diverse features are dynamically integrated via a two-stage fusion process: cross-modal attention followed by the iterative NAF block, which learns an adaptive fusion algorithm. Intrinsic explainability is embedded through path-specific spatial attention, Vim $\Delta$-value maps, traditional feature SE-attention, and cross-modal attention weights. Experiments on a diverse 9-class multi-organ medical image dataset demonstrate EVM-Fusion’s strong classification performance, achieving 99.75% test accuracy and provide multi-faceted insights into its decision-making process, highlighting its potential for trustworthy AI in medical diagnostics.
keywords:
@@ 1 Introduction
The accurate and timely classification of medical images is paramount in clinical diagnostics and treatment planning, spanning diverse modalities such as X-ray, Computed Tomography (CT), and Magnetic Resonance Imaging (MRI). While human expertise remains crucial, the increasing volume and complexity of medical imaging data presents significant challenges, including inter-observer variability and diagnostic delays [1]. Artificial Intelligence (AI), particularly deep learning, has demonstrated remarkable potential in automating and augmenting medical image analysis, promising to enhance both diagnostic efficiency and accuracy [2, 3, 4].
Early AI approaches often relied on handcrafted features capturing predefined image properties (e.g., texture, shape) [5, 6], which were then fed into traditional machine learning classifiers [7, 8, 9]. With the widespread application of deep learning technology, especially convolutional neural networks (CNNs) [3], CNN-based approaches often focus either on direct classification using learned features [10, 11, 12, 13, 14] or leverage image segmentation [15] to isolate regions of interest (ROIs) before classification [16, 17, 18, 19]. Despite their success, these models often operate as ”black boxes,” limiting clinical trust and adoption. Furthermore, fusing information from different architectural paradigms or feature types to achieve superior performance and generalizability remains a complex challenge. Simple concatenation or weighted averaging of features may not capture the intricate, context-dependent relationships between different feature streams.
To address these limitations, this paper proposes an Explainable Vision Mamba Architecture with Neural Algorithmic Fusion (EVM-Fusion), a novel deep learning architecture designed for robust and interpretable multi-organ medical image classification. It is a novel multi-path architecture integrating Vision Mamba (Vim) [20] for enhanced feature representation from diverse sources. What’s more, it introduces of the Neural Algorithmic Fusion (NAF) block, a new iterative and adaptive fusion mechanism. Finally, the comprehensive integration of intrinsic XAI components [21] on the model provides multi-level interpretability for medical image classification. Experimental results on a challenging multi-organ medical image dataset (comprising X-rays and MRIs for 9 distinct classes) demonstrate the superior performance and robustness of EVM-Fusion, alongside providing rich insights through its explainability features. The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 details the EVM-Fusion architecture. Section 4 describes the experimental setup. Section 5 presents and discusses the results, including quantitative performance and qualitative XAI analysis. Section 6 concludes the paper.
@@ 2 Related Work
2.1 Deep Learning Architectures for Medical Image Classification
Convolutional neural networks (CNNs) have long been the dominant paradigm in medical image analysis [3]. Pioneering architectures such as AlexNet [22], VGG [23], ResNet [24], and DenseNet [25] have established powerful hierarchical feature learning capabilities. ResNet solves the vanishing gradient problem in deep networks through residual connections, while DenseNet promotes feature reuse and reduces parameters through dense connections between layers. Dual Path Networks (DPNs) [26] aim to combine the advantages of ResNet and DenseNet. U-Net [27] was originally designed for biomedical image segmentation. Due to its effective multi-scale feature extraction and local detail representation capabilities, it has also been used for classification tasks [18], and sometimes DenseNet modules are integrated in its encoder, mainly for segmentation [28, 29, 30, 31]. Recently, Vision Transformers (ViT) [32] and its variants, such as the Swin Transformer [33], have shown promise by applying the Transformer [34] architecture, which originally achieved success in natural language processing (NLP), to vision tasks and capturing global context through self-attention mechanisms.
2.2 Multi-Path and Multi-Modal Fusion Techniques
Integrating information from multiple sources or pathways is a common strategy to improve model performance. In medical imaging, this may involve fusing features from different network architectures (multi-path) or different imaging modalities (multi-modality). While simple concatenation or averaging, while effective, may not optimally capture complex inter-modality dependencies. More advanced fusion techniques involve attention mechanisms that dynamically weigh the contributions of different feature sets. Learning an explicit adaptive fusion algorithm represents a step towards more sophisticated integrations.
@@ 2.2 Multi-Path and Multi-Modal Fusion Techniques
2.3 Explainable AI (XAI) in Medical Imaging
The ”black box” nature of many deep learning models hinders their clinical adoption, making XAI crucial [35]. Common post-hoc XAI methods include LIME [36] and SHAP [37]. For CNNs, activation mapping techniques like CAM [38] and its derivatives (Grad-CAM, Grad-CAM++) [39] highlight image regions influencing predictions. However, building intrinsically interpretable models, where explainability is a core part of architecture, is an increasingly important research direction [40]. My EVM-Fusion incorporates multiple intrinsic XAI components at different stages.
2.4 State Space Models (SSMs) and Mamba in Vision
In recent years, state-space models (SSMs) have emerged as a powerful alternative to Transformer sequence modeling, offering linear time complexity and strong long-range dependency capture capabilities. The Mamba architecture [41] introduced a selective scanning mechanism (S6 model) to achieve input-dependent state transition modulation. This has driven the rapid development of Mamba in vision tasks. Vim and the visual state-space model (VMamba) [42] apply the advantages of Mamba to 2D image data by splicing images into sequences and adopting specialized scanning strategies (e.g., bidirectional, multidirectional). These models show comparable performance to CNNs and ViT, while being more efficient in processing long sequences.
2.5 My Contribution in Context
While existing works have explored multi-path CNNs, attention-based fusion, and the application of Mamba to vision, EVM-Fusion introduces several novel contributions. Unlike standard multi-path models that often use simple fusion, EVM-Fusion pioneers a Neural Algorithmic Fusion (NAF) block, which learns an iterative, adaptive fusion process guided by a controller, following an initial cross-modal attention stage. This provides a more flexible and potentially more powerful fusion paradigm than fixed attention or concatenation schemes. Furthermore, EVM-Fusion integrates Vision Mamba modules into distinct feature extraction pathways (DenseNet-based and U-Net-based) to leverage their long-range modeling capabilities for different types of feature hierarchies, alongside a traditional feature path. Critically, it embeds multiple intrinsic XAI mechanisms across these paths and the fusion module, offering a more holistic and multi-faceted interpretability than typically found in complex fusion models. This combination of Mamba-enhanced diverse feature streams, a novel algorithmic fusion approach, and deep intrinsic explainability for multi-organ medical image classification distinguishes EVM-Fusion from prior art.
@@ 3 Model Architecture
The overall framework, illustrated in Fig. 1, is a multi-path network that integrates deep learning features from diverse backbone architectures with traditional handcrafted features, all enhanced by state-of-the-art sequence modeling and attention mechanisms. A key novelty lies in its two-stage fusion strategy, culminating in a Neural Algorithmic Fusion (NAF) block that learns an iterative process to combine information. Intrinsic explainability is a core design principle, with XAI components embedded throughout the architecture.
@@ 3.1 Overall Framework
Given an input medical image $I\in\mathbb{R}^{H\times W\times C}$ (where $H,W,C$ are height, width, and channels respectively), EVM-Fusion employs three parallel pathways to extract a rich set of complementary features: A DenseNet-Mamba path, designed to capture global contextual information and intricate feature patterns. A U-Net-Mamba path, focused on extracting multi-scale spatial hierarchies and detailed local information. A Traditional Features-Attention path, providing explicit, domain-relevant texture and statistical features.
Each path $p$ in those three produces a feature vector $v_{p}\in\mathbb{R}^{d_{model\_fusion}}$. These vectors are then fed into the Core Fusion Module, which first applies Cross-Modal Attention to model inter-path dependencies, followed by the Neural Algorithmic Fusion (NAF) block to generate a final, unified feature representation $v_{fused}\in\mathbb{R}^{d_{model\_fusion}}$. This $v_{fused}$ is then passed to a classification head to predict the image class.
@@ 3.2 Feature Extraction and Enhancement Paths
Each deep learning path (DenseNet and U-Net) incorporates a Vision Mamba module for advanced sequence-based feature refinement and a path-specific spatial attention mechanism for feature recalibration and explainability.
@@ 3.2.1 DenseNet-Mamba Path
This path aims to extract robust global and hierarchical features. The input image $I$ is first processed by a pre-trained DenseNet-121 backbone (excluding its final classification layer) to obtain a rich feature map $F_{D}=DenseNet(I)\in\mathbb{R}^{H_{D}\times W_{D}\times C_{D}}$. The feature map is then fed into a Vision Mamba Module. This module first uses a convolutional patch embedding layer to transform $F_{D}$ into a sequence of patch tokens $X_{D}\in\mathbb{R}^{N_{p}\times d_{mamba}}$, where $N_{p}$ is the number of patches and $d_{mamba}$ is the Mamba model dimension. The core of this module consists of $L_{M}$ stacked Mamba (S6) blocks [41] which process the sequence bidirectionally. Each Mamba block performs:
$x_{k}^{\prime},z_{k}=W_{in}x_{k}+b_{in}$ (1)
$x_{k}^{\prime\prime}=Conv1D(SiLU(x_{k}^{\prime}))$ (2)
$h_{k}=\bar{A}_{k}h_{k-1}+\bar{B}_{k}x_{k}^{\prime\prime}$ (3)
$y_{k}^{ssm}=C_{k}h_{k}$ (4)
$y_{k}=W_{out}(SiLU(z_{k})\odot y_{k}^{ssm}+Dx_{k}^{\prime\prime})$ (5)
for the $k$-th patch $x_{k}$. Where $\bar{A}_{k}=exp(\Delta_{k}A)$ and $\bar{B}_{k}\approx exp(\Delta_{k}B_{k})$ (Euler approximation of the continuous system $h^{\prime}(t)=Ah(t)+Bx(t)$). The parameter $\Delta_{k}$ acts as an input-dependent gate controlling the influence of $x_{k}^{\prime\prime}$ versus $h_{k-1}$, and its values are stored for XAI.
Thus, the output of the Vim module is to reshape $F_{D}$ into a feature map $F_{D,map}^{\prime}\in\mathbb{R}^{H_{D}^{\prime}\times W_{D}^{\prime}\times d_{model\_fusion}}$. This feature map is then processed by the spatial attention module, which generates a two-dimensional attention map.
$\begin{split}M_{spat}(F_{D,map}^{\prime})=\sigma(W_{conv}([AvgPool_{ch}(F_{D,map}^{\prime});\\ MaxPool_{ch}(F_{D,map}^{\prime})]))\end{split}$ (6)
The final feature vector for this path is $v_{D}=GAP(F_{D,map}^{\prime}\odot M_{spat}(F_{D,map}^{\prime}))\in\mathbb{R}^{d_{model\_fusion}}$. Where $GAP()$ means Global Average Pooling and the map $M_{spat}(F_{D,map}^{\prime})$ is saved as the Spatial Attn Map I in Fig. 1.
@@ 3.2.2 U-Net-Mamba Path
This path is designed to capture multi-scale spatial information and fine-grained details critical for medical image analysis. And its processing process is similar to Section 3.2.1.
@@ 3.2.3 Traditional Features-Attention Path
This path explicitly incorporates domain-relevant handcrafted features. For a grayscale version of input $I$, Gray-Level Co-occurrence Matrix (GLCM) features (e.g., contrast, homogeneity, energy over multiple distances and angles) and Local Binary Pattern (LBP) histogram features are extracted, resulting in a raw feature vector $f_{raw}\in\mathbb{R}^{d_{raw}}$.
Since traditional features are not very likely to be meaningfully organized into sequences, they are not suitable for processing using the Mamba module. Instead, I applied an multi-head attention layer [34] to learn the importance of different feature dimensions.
$f_{trad}^{\prime}=MHA(f_{raw})\in\mathbb{R}^{d_{raw}}$ (7)
It is then passed through an SEBlock [43]
$s=\sigma(W_{2}ReLU(W_{1}f_{trad}^{\prime}))$ (8)
Where $W_{1}\in\mathbb{R}^{(d_{raw}/r)\times d_{raw}}$ and $W_{2}\in\mathbb{R}^{d_{raw}\times(d_{raw}/r)}$ are linear layers, $r$ is the reduction ratio. The output $f_{trad}^{\prime\prime}=s\odot f_{trad}^{\prime}$ and the channel-wise attention scores $s\in\mathbb{R}^{d_{raw}}$ are saved as Trad SE Channel Attns for XAI. Finally, $f_{trad}^{\prime\prime}$ is passed through a fully connected layer with ReLU activation and Batch Normalization to produce $v_{T}\in\mathbb{R}^{d_{model\_fusion}}$
Figure 2: The Confusion Matrices of the 10 models evaluated on the test set. To demonstrate the rationality of the model architecture, I conducted ablation experiments. In the figure, EVM-DUHF represents the complete EVM-Fusion model with the DenseNet-Mamba Path, U-Net-Mamba Path, Traditional Features-Attention Path. The other three models in the figure, EVM-DHF, EVM-DU, EVM-UHF, are variants of EVM-DUHF with the U-Net-Mamba Path, Traditional Features-Attention Path, DenseNet-Mamba Path removed, respectively. Additionally, U-DenseNet refers to the model described in [28], which uses DenseNet-121 as the encoder for U-Net.

### v4

@@ Abstract
Medical image classification is critical for clinical decision-making, yet demands for accuracy, interpretability, and generalizability remain challenging. This paper introduces EVM-Fusion, an Explainable Vision Mamba architecture featuring a novel Neural Algorithmic Fusion (NAF) mechanism for multi-organ medical image classification. EVM-Fusion leverages a multipath design, where DenseNet and U-Net based pathways, enhanced by Vision Mamba (Vim) modules, operate in parallel with a traditional feature pathway. These diverse features are dynamically integrated via a two-stage fusion process: cross-modal attention followed by the iterative NAF block, which learns an adaptive fusion algorithm. Intrinsic explainability is embedded through path-specific spatial attention, Vim $\Delta$ -value maps, traditional feature SE-attention, and cross-modal attention weights. Experiments on a diverse 9-class multi-organ medical image dataset demonstrate EVM-Fusion’s strong classification performance and provide multi-faceted insights into its decision-making process, highlighting its potential for trustworthy AI in medical diagnostics.
{IEEEkeywords}
@@ 1 Introduction
[…]
The accurate analysis of medical images is fundamental to modern clinical practice [1, 2, 3, 4]. A standard and effective workflow for detailed diagnosis involves two stages. First, a model segments [5] a specific region of interest (ROI). Then, a second model classifies that ROI. This segmentation-based approach is precise [6, 7, 8]. However, it requires detailed pixel-level annotations, which are often expensive and time-consuming to create.
This paper explores a different but equally important clinical problem: automated image triage. A triage system needs to make a rapid, holistic assessment of an entire medical image. Its goal is to quickly categorize the image, such as identifying a potential abnormality or its likely type. This can help clinicians prioritize urgent cases and manage workflows more efficiently, especially in settings with a high volume of patients.
Building an effective end-to-end model for this triage task is challenging. The model must analyze the whole image without explicit guidance. This creates three specific problems. First, the model needs to understand both fine-grained local details and long-range global context. For example, it must see both the texture of a small lesion and its position relative to the entire organ. Standard Convolutional Neural Networks (CNNs) [3, 9, 10] are good at local feature extraction but often struggle to capture global context efficiently. Second, a single image contains very different kinds of information. Deep learning models extract powerful hierarchical features. At the same time, traditional handcrafted features can provide robust and explicit information about texture or statistics [11, 12, 13, 14, 15]. An effective triage model should use all these heterogeneous feature types. However, fusing them is a difficult problem. Simple methods like feature concatenation often fail because they cannot model the complex, non-linear relationships between these different feature streams. Third, many complex deep learning models operate like a ”black box”. This lack of transparency is a major barrier to trust and adoption in clinical settings. A reliable triage system must be interpretable, allowing clinicians to understand the basis of its decisions.
To solve these specific challenges, we designed and validated the Explainable Vision Mamba Architecture with Neural Algorithmic Fusion (EVM-Fusion). Our work is not about creating a single new component. Instead, our contribution is a synergistic framework that intelligently integrates multiple advanced components. Each component is chosen to address one of the problems mentioned above. We propose: A Mamba-enhanced multi-path architecture to capture heterogeneous features. The framework has three paths. A DenseNet path extracts robust semantic features. A U-Net path focuses on multi-scale details. A third path provides traditional texture features. We enhance the two deep learning paths with Vision Mamba (Vim) [16] modules to specifically address the challenge of modeling long-range dependencies. A learnable, two-stage fusion strategy to integrate these diverse features. A Cross-Modal Attention stage first lets the different paths exchange contextual information. Then, a Neural Algorithmic Fusion (NAF) block learns a multi-step, iterative process to find the best way to combine the features. This design directly addresses the limitations of simpler, fixed fusion rules. A holistically integrated explainability (XAI) system. XAI [17] mechanisms are built into every stage of the model. This provides a multi-faceted view of the model’s reasoning process, which is essential for building a trustworthy AI tool for clinicians.
We tested our framework on a challenging 9-class dataset. Our experiments show that the complete EVM-Fusion system outperforms powerful state-of-the-art models from different architectural families. Our ablation studies further prove that every component in our synergistic design is necessary for achieving this top performance. The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 details the EVM-Fusion architecture. Section 4 describes the experimental setup. Section 5 presents and discusses the results, and Section 6 concludes the paper.
@@ 2 Related Work
The field of medical image analysis has two main paradigms. The first is a two-stage, segmentation-based approach. Models like U-Net [18] are often used to delineate specific ROIs [19, 20, 21, 22], which are then classified. This method is the standard for precise clinical diagnosis. The second paradigm is holistic, end-to-end classification [23, 24, 25, 26, 27]. Here, a model processes the entire image directly to make a prediction. Our work contributes to this second paradigm. We aim to develop a robust framework for rapid image triage, a task where holistic analysis can be more practical than detailed segmentation.
2.1 Architectures for Visual Representation
CNNs like ResNet [28] and DenseNet [29] have been foundational in medical imaging. They are very effective at learning local patterns from images. However, their fixed and local receptive fields are a known limitation for modeling global context. Vision Transformers (ViT) [30] and Swin Transformer [31] address this problem using a self-attention mechanism from natural language processing. More recently, State Space Models (SSMs) have emerged as another powerful tool for sequence modeling. Mamba [32] is a key example. Vision-specific versions like VMamba [33] and Vim adapt this technology for images. They can capture long-range dependencies with high efficiency. Our framework recognizes the distinct strengths of these architectures. We use CNNs for their proven ability to extract local features and enhance them with Mamba to add the crucial capability of global context modeling.
2.2 Feature Fusion in Deep Learning
Combining features from multiple sources is a common strategy to improve model performance. Simple approaches include feature concatenation or element-wise averaging. These methods are straightforward but may not be optimal [34]. More advanced techniques use attention mechanisms to dynamically assign weights to different features [35]. However, these methods typically use a single-step weighting scheme. This can be too rigid for fusing features that are highly diverse, such as combining deep hierarchical features with traditional statistical features. This specific challenge motivates our use of Neural Algorithmic Fusion (NAF) [bib21]. NAF learns a more complex, multi-step fusion process, which is better suited for this difficult task.
2.3 Explainable AI (XAI) in Medical Imaging
The ”black box” nature of many deep learning models hinders their clinical adoption, making XAI crucial [36]. Common post-hoc XAI methods include LIME [37] and SHAP [38]. For CNNs, activation mapping techniques like CAM [39] and its derivatives (Grad-CAM, Grad-CAM++) [40] highlight image regions influencing predictions. However, building intrinsically interpretable models, where explainability is a core part of the architecture, is an increasingly important research direction [41]. Our EVM-Fusion incorporates multiple intrinsic XAI components at different stages.
@@ 3 Model Architecture
The EVM-Fusion architecture is shown in Fig 1. It is a multi-path network. We designed it to address the challenges of holistic image triage. A key part of our design is the two-stage fusion strategy. This strategy uses a Neural Algorithmic Fusion (NAF) block to learn an iterative process for combining information. The architecture also has intrinsic XAI components to ensure transparency.
@@ 3.1 Overall Framework
Given an input medical image $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ , EVM-Fusion employs three parallel pathways to extract a rich set of complementary features. The motivation for this multi-path design is to capture different facets of the image data simultaneously: A DenseNet-Mamba path: Designed to capture robust, high-level semantic features by leveraging the deep feature reuse of DenseNet. A U-Net-Mamba path: Focused on extracting multi-scale spatial hierarchies and preserving fine-grained local details, a strength of the U-Net architecture’s encoder-decoder structure. A Traditional Features-Attention path: Included to provide explicit, domain-relevant texture and statistical features that deep models might overlook.
Each path $p$ produces a feature vector $\mathbf{v}_{p}\in\mathbb{R}^{d_{\text{model\_fusion}}}$. These vectors are then passed to the Core Fusion Module, which first applies Cross-Modal Attention to model inter-path dependencies, followed by the NAF block to generate a final, unified feature representation $\mathbf{v}_{\text{fused}}\in\mathbb{R}^{d_{\text{model\_fusion}}}$. This vector is then used by a classification head to predict the image class.
@@ 3.2 Feature Extraction and Enhancement Paths
Each deep learning path incorporates a Vision Mamba module for advanced sequence-based feature refinement and a path-specific spatial attention mechanism.
@@ 3.2.1 DenseNet-Mamba Path
This path aims to extract robust semantic features. The input image $\mathbf{I}$ is first processed by a pre-trained DenseNet-121 backbone to obtain a feature map $\mathbf{F}_{D}=DenseNet(\mathbf{I})\in\mathbb{R}^{H_{D}\times W_{D}\times C_{D}}$. To model long-range dependencies, this feature map is processed by a Vision Mamba (Vim) module. It first transforms $\mathbf{F}_{D}$ into a sequence of patch tokens $\mathbf{X}_{D}\in\mathbb{R}^{N_{p}\times d_{mamba}}$. The core of the module consists of stacked Mamba (S6) blocks [32] that process the sequence. In brief, for the $k$-th patch token $\mathbf{x}_{k}$, a Mamba block computes an output $\mathbf{y}_{k}$ via a selective state space model. The key state update is:
$\mathbf{h}_{k}=\bar{\mathbf{A}}_{k}\mathbf{h}_{k-1}+\bar{\mathbf{B}}_{k}\mathbf{x}_{k}^{\prime\prime}$ (1)
Where $\mathbf{x}_{k}^{\prime\prime}$ is a projection of the input $\mathbf{x}_{k}$. The matrices $\bar{\mathbf{A}}_{k}$ and $\bar{\mathbf{B}}_{k}$ are parameterized by a data-dependent scalar $\mathbf{\Delta}_{k}$, which acts as a selective gate. This gate modulates the flow of information, allowing the model to either focus on the current input or maintain its previous state, effectively learning long-range dependencies. The $\mathbf{\Delta}_{k}$ values are stored for XAI analysis. The Vim module’s output is reshaped back into a feature map $\mathbf{F}_{D,map}^{\prime}\in\mathbb{R}^{H_{D}^{\prime}\times W_{D}^{\prime}\times d_{\text{model\_fusion}}}$.
To recalibrate the feature map spatially, a spatial attention module is applied:
$\begin{split}\mathbf{M}_{\text{spat}}(\mathbf{F^{\prime}})=\sigma(\mathbf{W}_{\text{conv}}[\text{AvgPool}_{\text{ch}}(\mathbf{F^{\prime}});\\ \text{MaxPool}_{\text{ch}}(\mathbf{F^{\prime}})])\end{split}$ (2)
where channel-wise average and max pooling aggregate feature information, and $\sigma$ is the sigmoid function. The final path vector is $\mathbf{v}_{D}=GAP(\mathbf{F}^{\prime}\odot\mathbf{M}_{\text{spat}}(\mathbf{F}^{\prime}))$, where GAP is Global Average Pooling. The attention map $\mathbf{M}_{\text{spat}}$ is saved for XAI.
@@ 3.2.2 U-Net-Mamba Path
This path is designed to capture multi-scale spatial information and fine-grained details, which are often critical for medical image analysis. It employs a full U-Net architecture as the initial feature extractor. The output from the U-Net’s final decoder layer, a feature map $\mathbf{F}_{U}$, is then processed in a manner identical to the DenseNet path: it is passed through its own dedicated Vision Mamba and Spatial Attention modules to produce the feature vector $\mathbf{v}_{D}\in\mathbb{R}^{d_{\text{model\_fusion}}}$.
@@ 3.2.3 Traditional Features-Attention Path
This path explicitly incorporates domain-relevant handcrafted features. For a grayscale version of input $\mathbf{I}$, Gray-Level Co-occurrence Matrix (GLCM) and Local Binary Pattern (LBP) features are extracted into a raw feature vector $\mathbf{f}_{raw}\in\mathbb{R}^{d_{raw}}$.
As these features lack a natural sequential order, a Multi-Head Attention (MHA) layer [42] is used instead of Mamba. Here, MHA learns inter-feature relationships, allowing features to be re-weighted based on the context of other features in the vector.
$\mathbf{f}^{\prime}_{\text{trad}}=\text{MHA}(\mathbf{f}_{\text{raw}},\mathbf{f}_{\text{raw}},\mathbf{f}_{\text{raw}})$ (3)
This is followed by a Squeeze-and-Excitation (SE) block [43] to perform feature-wise recalibration:
$\mathbf{s}=\sigma(\mathbf{W}_{2}\text{ReLU}(\mathbf{W}_{1}\mathbf{f}^{\prime}_{\text{trad}}))$ (4)
Where $\mathbf{W}_{1}\in\mathbb{R}^{(d_{raw}/r)\times d_{raw}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{raw}\times(d_{raw}/r)}$ are linear layers, $r$ is the reduction ratio. The output $\mathbf{f}_{trad}^{\prime\prime}=\mathbf{s}\odot\mathbf{f}_{trad}^{\prime}$ , and the scores s are saved for XAI. Finally, a fully connected layer produces $\mathbf{v}_{T}\in\mathbb{R}^{d_{\text{model\_fusion}}}$.